# Enhancing Consistency-Based Image Generation via Adversarialy-Trained Classification and Energy-Based Discrimination

### v1

@@ 1 Introduction
Diffusion models, also known as score-based generative models,  [13],  [24],  [27],  [28], [29] have set new benchmarks in multiple fields, the prime of which is image generation  [4],  [22]. A central aspect of these models is the iterative sampling process that gradually eliminates noise from an initial random vector, this way carving fair samples from a prior distribution. By controlling the depth of such processes, these methods offer a flexible balance between computational effort and sample quality. However, compared to single-step generative models like GANs  [9], the iterative generation procedure of diffusion models typically requires 10–2000 times more compute for sample generation  [29],  [28],  [13], causing slow inference and limited real-time applications. Addressing this shortcoming, Consistency models  [26] have been developed to expedite the sampling process, offering a faster alternative with one or two-step data generation. Unfortunately, their reliance on an intricate distillation process tends to limit the ultimate quality of the generated images.
Recent work that aims to remedy this performance gap integrates Consistency Models  [26] with Generative Adversarial Networks, so as to tackle common GAN issues like mode collapse and unstable training process, thereby enhancing the diversity and stability of generated images [20]. Adversarial Consistency Training (ACT)  [17] further refines this approach by incorporating a discriminator within the Consistency training framework, improving the generative model’s quality by fine-tuning the image generation process for more realistic outputs. These methods offer new ways to overcome limitations in image fidelity, but require long and computationally expensive training.
@@
Past research has shown that deep neural networks can be easily deceived by adversarial attacks  [10],  [18],  [30]. Therefore, numerous strategies for making these networks more robust were developed, with adversarial training becoming one of the most favored approaches  [21],  [10]. During the analysis of such classifiers, a notable phenomenon was discovered: perceptually aligned gradients  [31],  [5]. This trait implies that adjustments to an image, guided by the robust classifier, result in changes that are visually consistent with the intended classification.
Our method utilizes the above phenomenon and iteratively modifies the images created by Consistency-models using a joint robust classifier-discriminator. While the classifier aims to maximize the conditional probability of a target class, the discriminator portion of the very same network uses the softmax values to maximize the energy difference between synthesized and observed data, functioning as an Energy-based Model  [19].
@@
    We show a significant quantitative and qualitative improvements in the perceptual quality of synthesized images for both Consistency-Training and Consistency-distillation, with an FID boost of $27.48\%$ and $20.96\%$, respectively.
@@
    We demonstrate a remarkable capability of adversarial robust classifiers: Serving as a joint classifier-discriminator informed of both synthesized and observed data distributions. This fusion leads to enhanced effectiveness of projected-gradient steps, leading to an additional improvement of $11.2\%$ in FID results, over a classifier boosting alone  [6].
@@ 2.1 Diffusion Models and Consistency
Diffusion models  [29],  [14] generate data (e.g., images) by starting from a plain random vector and manipulating it progressively by sequential denoising steps and noise perturbations. More specifically, let $p_{0}(x)$ denote the Probability Density Function (PDF) of true images. Diffusion models introduce a forward path, in which $p_{0}(x)$ is gradually diffused over the time interval $0\leq t\leq T$ to a canonical Gaussian, $p_{T}(x)={\cal N}(0,I)$. This is formulated via a Stochastic Differential Equation (SDE) that describes a continuous noise-mixing process,
@@
Diffusion models are all about designing the reverse process from $t=T$ to $t=0$, in which noise is filtered gradually to produce high quality images that serve as fair samples from $p_{0}(x)$. While the original diffusion techniques  [13] offer an SDE-based reverse path, the more recent and more effective alternatives (e.g. DDIM [25]) remove the intermediate stochasticity and offer an ODE-based reverse process of the form
@@
In practice, Diffusion Models are an Euler-Maruyama discretization [16] of the above equation, resulting with a chain of iterations that starts with a random canonical Gaussian noise $x_{T}$, and gradually moves $x_{t}$ towards $x_{0}$ by steps of denoising that follow the above equation. The resulting $x_{0}$ can be considered as a fair sample from $p_{0}(x)$, assuming that the denoiser is well-trained.
Consistency models  [26] are heavily inspired by the theory described above, aiming to preserve the end-to-end behavior that takes us from $x_{T}$ to $x_{0}$, but while expediting the process to a single (or a few) step. Given a solution trajectory $\{x_{t}\}_{t\in[0,T]}$ of the above reverse equation, the Consistency function is the trained network $f:(x_{t},t)\rightarrow x_{0}$. This function should have the property of self-consistency: For two temporal points $t,s\in[0,T]$ and their corresponding stages in the same trajectory, $x_{t}$ and $x_{s}$, the outputs should be the same $f(x_{t},t)=f(x_{s},s)=x_{0}$, thus explaining the name “Consistency”. Such models can be trained in two main ways  . The first method (CD) involves distilling knowledge from a pre-trained diffusion model, allowing it to learn the distribution directly from the more computationally intensive yet well-performing diffusion process. The second method (CT) trains the function $f(\cdot,\cdot)$ directly via a tailored loss, without the need for a pre-existing diffusion model.
@@ 2.2 Boosting Synthesis via Robust Classifier (BIGROC)
Consider a goal of refining generated images, so as to improve their perceptual quality. BIGROC  [6] offers such a post-processing mechanism, by leveraging a robust classifier. Let $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ be a classifier that maps images in $\mathbb{R}^{d}$ to scores for $C$ classes, also known as logits, and denote it’s $c$th output by $f_{\theta}^{c}(x)$. These logits aim to have a probabilistic interpretation, being related to the estimated posterior probability $p_{\theta}(c|x)$. Starting with an arbitrary image $x$ and taking gradient steps $\nabla_{x}f_{\theta}^{c}(x)$ in order to maximize $f_{\theta}^{c}(x)$, one would expect this process to produce an altered image that better resembles the targeted class $c$. However, when using a plain classifier, even if it is very well performing, its gradients behave as a perceptually meaningless noise that is unrelated to the image content. This implies that, while the resulting image would tend to be classified to $c$, its visual appearance would be hardly changed from the original image $x$. This phenomenon is well-known, giving rise to the wide topic of adversarial attacks on neural networks  [21],  [9].
@@
Past research has shown that, as opposed to the above, robust classifiers produce Perceptually Aligned Gradients (PAG) [31, 15, 1, 8, 7]. These gradients are a trait of adversarially trained models, their content being consistent with the human visual perception  [31],  [5]. This implies that when used within the gradient steps mentioned above that aim to pull $x$ towards class $c$, robust models are expected to yield meaningful features aligned with the target class, thus modifying $x$ in a perceptually convincing way.
@@
Harnessing the above knowledge, the BIGROC algorithm  [6] proposes an iterative process for boosting the perceptual quality of synthesized images, assuming that they are classifiable via $f_{\theta}(x)$. This is obtained by modifying any generated image $x$ so as to maximize the posterior probability of a given target class $\hat{c}$, i.e, $p_{\theta}(\hat{c}|x)$, where $p_{\theta}$ is modeled by an adversarially trained classifier and ${\hat{c}}$ stands for the estimated class the image belongs to. The outcome of this process is a modified version of $x$ that is more aligned with the class ${\hat{c}}$ from a human point of view, thus having a better perceptual quality.
@@
While there are various techniques for generating adversarial examples, the one used in BIGROC is Targeted Projected Gradient Descent  [21] – a deterministic iterative process that operates as described in  1 $\Pi_{\varepsilon}$ in this algorithm stands for the projection operator that ensures that the resulting image is $\varepsilon$-close to the starting image, and $\mathcal{\ell}$ is the loss function (e.g., cross-entropy) that defines how the distance between the classifier output and the target label is computed. For simplicity, we can omit the loss and use the gradient of the softmax output of the classifier directly, $\delta_{t+1}=\Pi_{\varepsilon}(\delta_{t}-\nabla_{\delta}f_{\theta}^{\hat{c}}(x+\delta_{t}))$.
@@
We would like to draw the readers attention to the fact that BIGROC  [6] is totally unaware of the generated images’ distribution, which it aims to improve. This brings us naturally towards the proposed strategy in this paper.
@@ 2.3 Energy-Based Models
The core idea behind Energy-Based Models (EBM)  [19] is turning an arbitrary function into a probability distribution. Consider a learned energy function $E_{\theta}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$, where $x\in\mathbb{R}^{d}$ is an input image, and it’s output is a scalar value in the range (,+)($\infty$,$\infty$)( - ∞ , + ∞ ). This neural network induces a PDF $p_{\phi}(x)$ that has a Gibbs form:
@@
EBMs are a family of algorithms that aim to learn $E_{\theta}(x)$ such that the resulting $p_{\theta}(x)$ aligns with the probability density function of the given training set. Given a learned EBM, image synthesis can be obtained by iteratively climbing the function $p_{\theta}(x)$ using Stochastic Gradient Langevin Dynamics (SGLD) [32]. PGD, as described above, offers an appealing deterministic alternative to SGLD, by leveraging a loss-function and an additional $\varepsilon$-ball constraint in it’s computation.
Energy-based models and image classifiers can be shown to have common grounds  [11], as detailed in Appendix  A. Assuming we have a classifier $f_{\theta}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ that maps images to scores for $C$ classes, the probability of the $c^{th}$ label is represented using the Softmax function,
@@ 3 Proposed Approach
In this work we propose a method to enhance the perceptual quality of images generated by Consistency-based techniques  [26]. Our approach utilizes a robust joint classifier-discriminator to guide synthesized images to better resemble real images. In section 3.1, we explore this method’s training objectives and technical details. The post-processing procedure, described in section 3.2, uses the guidance of the joint model in an iterative algorithm to refine the realism and visual appeal of synthesized images. In this process, the synthesized images, originating from the generated distribution $p_{\theta}$, are modified to align with the real images distribution $p_{data}$.
@@ 3.1 Training a Joint Classifier-Discriminator
Our method aims to leverage the Perceptually Aligned Gradients (PAG) property  [31], and further extend and improve refinement capabilities of models that possess this trait by introducing the model with the distribution of both $p_{\theta}$ and $p_{data}$. To this end, we empower the classifier model by incorporating a robust discriminator within it, so as to improve guidance performance significantly with this joint model. This dual-training approach allows the proposed model to better differentiate and understand the unique features of both real and synthetic images.
@@ 3.1.1 Proposed Training objective
While Energy-Based Models (EBMs)  [19] may align perfectly with our task of improving image quality by modifying out-of-distribution samples to align with the real data manifold, several challenges impact their practical application. These include the sensitivity of EBM training to the choice of the energy function used, leading often to training instability and poor generalization performance. EBMs necessarily rely on computationally complex sampling techniques (e.g. MCMC), which add to their overall complexity and instability. More on these matters is brought in Appendix  B.
@@
The overall loss is simply an addition of the above two expressions, $\mathcal{L}(\theta)=\mathcal{L}_{CE}(\theta)+\mathcal{L}_{BCE}(\theta)$. This formulation ensures that the model learns to distinguish effectively between real and generated images, as well as classifying them correctly, enhancing its ability to deal with adversarially altered images and to generalize well across different types of image manipulations. More importantly, due to the adversarial training, the learned model has meaningful gradients (i.e., PAG  [31],  [5]), which are about to play a vital role in our inference process of pushing Consistency generated images to their improved versions. As our training strategy considers both classification and discrimination losses, these gradients carry both improvements geared by class adjustments and a migration from $p_{\theta}$ to $p_{data}$.
In our training process, as described above, we initialize the joint model with a robust RN50 network ( [21]), while replacing its fully connected head with a randomly initialized one. As a consequence, we expect this network to inherit it’s original classification robustness, which proves to be valuable for the overall image improvement obtained hereafter. Our training stands on the shoulders of this classifier, as we update the model parameters with the loss described above using only attacked (real and synthesized) images.

### v2

# Enhancing Consistency-Based Image Generation via Adversarially-Trained Classification and Energy-Based Discrimination

@@ 1 Introduction
Diffusion models, also known as score-based generative models,  ho2020denoising ,  sohl2015deep ,  song2019generative ,  song2020improved , song2020score have set new benchmarks in multiple fields, the prime of which is image generation  dhariwal2021diffusion ,  nichol2021improved . A central aspect of these models is the iterative sampling process that gradually eliminates noise from an initial random vector, this way carving fair samples from a prior distribution. By controlling the depth of such processes, these methods offer a flexible balance between computational effort and sample quality. However, compared to single-step generative models like GANs  goodfellow2014generative , the iterative generation procedure of diffusion models typically requires 10–2000 times more compute for sample generation  song2020score ,  song2020improved ,  ho2020denoising , causing slow inference and limited real-time applications. Addressing this shortcoming, Consistency models  song2023consistency have been developed to expedite the sampling process, offering a faster alternative with one or two-step data generation. Unfortunately, their reliance on an intricate distillation process tends to limit the ultimate quality of the generated images.
Recent work that aims to remedy this performance gap integrates Consistency Models  song2023consistency with Generative Adversarial Networks, so as to tackle common GAN issues like mode collapse and unstable training process, thereby enhancing the diversity and stability of generated images lu2023cm . Adversarial Consistency Training (ACT)  kong2023act further refines this approach by incorporating a discriminator within the Consistency training framework, improving the generative model’s quality by fine-tuning the image generation process for more realistic outputs. These methods offer new ways to overcome limitations in image fidelity, but require long and computationally expensive training.
@@
The adversarial approach is critical in our training process, as we will further explain in section  3. We rely on past research that has shown that deep neural networks can be easily deceived by adversarial attacks  goodfellow2014explaining ,  kurakin2018adversarial ,  szegedy2014intriguing . Therefore, numerous strategies have been developed to make these networks more robust, with adversarial training becoming one of the most favored approaches  madry2018towards ,  goodfellow2014explaining . During the analysis of such classifiers, a notable phenomenon was discovered: perceptually aligned gradients  tsipras2018robustness ,  engstrom2019adversarial . This trait implies that adjustments to an image, guided by the robust classifier, result in changes that are visually consistent with the intended classification.
Our method utilizes the above phenomenon and iteratively modifies the images created by Consistency-models using a joint robust classifier-discriminator. While the classifier aims to maximize the conditional probability of a target class, the discriminator portion of the very same network uses the softmax values to maximize the energy difference between synthesized and observed data, functioning as an Energy-based Model  lecun2006tutorial .
@@
    We show significant quantitative and qualitative improvements in the perceptual quality of synthesized images for both Consistency-Training and Consistency-distillation, with an FID boost of $27.48\%$ and $20.96\%$, respectively.
@@
    We demonstrate a remarkable capability of adversarial robust classifiers: Serving as a joint classifier-discriminator informed of both synthesized and observed data distributions. This fusion leads to enhanced effectiveness of projected-gradient steps, leading to an additional improvement of $11.2\%$ in FID results, over a classifier boosting alone  ganz2022bigroc .
@@ 2.1 Diffusion Models and Consistency
Diffusion models  song2020score ,  karras2022elucidating generate data (e.g., images) by starting from a plain random vector and manipulating it progressively by sequential denoising steps and noise perturbations. More specifically, let $p_{0}(x)$ denote the Probability Density Function (PDF) of true images. Diffusion models introduce a forward path, in which $p_{0}(x)$ is gradually diffused over the time interval $0\leq t\leq T$ to a canonical Gaussian, $p_{T}(x)={\cal N}(0,I)$. This is formulated via a Stochastic Differential Equation (SDE) that describes a continuous noise-mixing process,
@@
Diffusion models are all about designing the reverse process from $t=T$ to $t=0$, in which noise is filtered gradually to produce high quality images that serve as fair samples from $p_{0}(x)$. While the original diffusion techniques  ho2020denoising offer an SDE-based reverse path, the more recent and more effective alternatives (e.g. DDIM song2020denoising ) remove the intermediate stochasticity and offer an ODE-based reverse process of the form
@@
In practice, Diffusion Models are an Euler-Maruyama discretization kloeden1992stochastic of the above equation, resulting with a chain of iterations that starts with a random canonical Gaussian noise $x_{T}$, and gradually moves $x_{t}$ towards $x_{0}$ by steps of denoising that follow the above equation. The resulting $x_{0}$ can be considered as a fair sample from $p_{0}(x)$, assuming that the denoiser is well-trained.
Consistency models  song2023consistency are heavily inspired by the theory described above, aiming to preserve the end-to-end behavior that takes us from $x_{T}$ to $x_{0}$, but while expediting the process to a single (or a few) step. Given a solution trajectory $\{x_{t}\}_{t\in[0,T]}$ of the above reverse equation, the Consistency function is the trained network $f:(x_{t},t)\rightarrow x_{0}$. This function should have the property of self-consistency: For two temporal points $t,s\in[0,T]$ and their corresponding stages in the same trajectory, $x_{t}$ and $x_{s}$, the outputs should be the same $f(x_{t},t)=f(x_{s},s)=x_{0}$, thus explaining the name “Consistency”. Such models can be trained in two main ways  song2023consistency . The first method (CD) involves distilling knowledge from a pre-trained diffusion model, allowing it to learn the distribution directly from the more computationally intensive yet well-performing diffusion process. The second method (CT) trains the function $f(\cdot,\cdot)$ directly via a tailored loss, without the need for a pre-existing diffusion model.
@@ 2.2 Boosting Synthesis via Robust Classifier (BIGROC)
Consider a goal of refining generated images, so as to improve their perceptual quality. BIGROC  ganz2022bigroc offers such a post-processing mechanism, by leveraging a robust classifier.
@@
Let $f_{\theta}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ be a classifier that maps images in $\mathbb{R}^{d}$ to scores for $C$ classes, also known as logits, and denote it’s $c$th output by $f_{\theta}^{c}(x)$. These logits aim to have a probabilistic interpretation, being related to the estimated posterior probability $p_{\theta}(c|x)$. Starting with an arbitrary image $x$ and taking gradient steps $\nabla_{x}f_{\theta}^{c}(x)$ in order to maximize $f_{\theta}^{c}(x)$, one would expect this process to produce an altered image that better resembles the targeted class $c$. However, when using a plain classifier, even if it is very well performing, its gradients behave as a perceptually meaningless noise that is unrelated to the image content. This implies that, while the resulting image would tend to be classified to $c$, its visual appearance would be hardly changed from the original image $x$. This phenomenon is well-known, giving rise to the wide topic of adversarial attacks on neural networks  madry2018towards ,  goodfellow2014generative .
@@
Past research has shown that, as opposed to the above, robust classifiers produce Perceptually Aligned Gradients (PAG) tsipras2018robustness ; kaur2019perceptually ; aggarwal2020benefits ; ganz2022perceptually ; Ganz_2024_WACV . These gradients are a trait of adversarially trained models, their content being consistent with the human visual perception  tsipras2018robustness ,  engstrom2019adversarial . This implies that when used within the gradient steps mentioned above that aim to pull $x$ towards class $c$, robust models are expected to yield meaningful features aligned with the target class, thus modifying $x$ in a perceptually convincing way.
@@
Harnessing the above knowledge, the BIGROC algorithm  ganz2022bigroc proposes an iterative process for boosting the perceptual quality of synthesized images, assuming that they are classifiable via $f_{\theta}(x)$. This is obtained by modifying any generated image $x$ so as to maximize the posterior probability of a given target class $\hat{c}$, i.e, $p_{\theta}(\hat{c}|x)$, where $p_{\theta}$ is modeled by an adversarially trained classifier and ${\hat{c}}$ stands for the estimated class the image belongs to. The outcome of this process is a modified version of $x$ that is more aligned with the class ${\hat{c}}$ from a human point of view, thus having a better perceptual quality.
@@
While there are various techniques for generating adversarial examples, the one used in BIGROC is Targeted Projected Gradient Descent  madry2018towards – a deterministic iterative process that operates as described in Algorithm  1, $\Pi_{\varepsilon}$ in this algorithm stands for the projection operator that ensures that the resulting image is $\varepsilon$-close to the starting image, and $\mathcal{\ell}$ is the loss function (e.g., cross-entropy) that defines how the distance between the classifier output and the target label is computed. For simplicity, we can omit the loss and use the gradient of the softmax output of the classifier directly,
$\delta_{t+1}=\Pi_{\varepsilon}(\delta_{t}-\nabla_{\delta}f_{\theta}^{\hat{c}}(x+\delta_{t})).$ (3)
@@
We would like to draw the readers attention to the fact that BIGROC  ganz2022bigroc is totally unaware of the generated images’ distribution, which it aims to improve. This brings us naturally towards the proposed strategy in this paper.
@@ 2.3 Energy-Based Models
The core idea behind Energy-Based Models (EBM)  lecun2006tutorial is turning an arbitrary function into a probability distribution. Consider a learned energy function $E_{\theta}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}$, where $x\in\mathbb{R}^{d}$ is an input image, and it’s output is a scalar value in the range (,+)($\infty$,$\infty$)( - ∞ , + ∞ ). This neural network induces a PDF $p_{\phi}(x)$ that has a Gibbs form:
@@
EBMs are a family of algorithms that aim to learn $E_{\theta}(x)$ such that the resulting $p_{\theta}(x)$ aligns with the probability density function of the given training set. Given a learned EBM, image synthesis can be obtained by iteratively climbing the function $p_{\theta}(x)$ using Stochastic Gradient Langevin Dynamics (SGLD) welling2011bayesian . PGD, as described above, offers an appealing deterministic alternative to SGLD, by leveraging a loss-function and an additional $\varepsilon$-ball constraint in it’s computation.
Energy-based models and image classifiers can be shown to have common grounds  grathwohl2019your , as detailed in Appendix  A. Assuming we have a classifier $f_{\theta}(x):\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ that maps images to scores for $C$ classes, the probability of the $c^{th}$ label is represented using the Softmax function,
@@ 3 Proposed Approach
In this work we propose a method to enhance the perceptual quality of images generated by Consistency-based techniques  song2023consistency . Our approach utilizes a robust joint classifier-discriminator to guide synthesized images to better resemble real images. In section 3.1, we explore this method’s training objectives and technical details. The post-processing procedure, described in section 3.2, uses the guidance of the joint model in an iterative algorithm to refine the realism and visual appeal of synthesized images. In this process, the synthesized images, originating from the generated distribution $p_{\theta}$, are modified to align with the real images distribution $p_{data}$.
@@ 3.1 Training a Joint Classifier-Discriminator
Our method aims to leverage the Perceptually Aligned Gradients (PAG) property  tsipras2018robustness , and further extend and improve refinement capabilities of models that possess this trait by introducing the model with the distribution of both $p_{\theta}$ and $p_{data}$. To this end, we empower the classifier model by incorporating a robust discriminator within it, so as to improve guidance performance significantly with this joint model. This dual-training approach allows the proposed model to better differentiate and understand the unique features of both real and synthetic images.
@@ 3.1.1 Proposed Training objective
While Energy-Based Models (EBMs)  lecun2006tutorial may align perfectly with our task of improving image quality by modifying out-of-distribution samples to align with the real data manifold, several challenges impact their practical application. These include the sensitivity of EBM training to the choice of the energy function used, leading often to training instability and poor generalization performance. EBMs necessarily rely on computationally complex sampling techniques (e.g. MCMC), which add to their overall complexity and instability. More on these matters is brought in Appendix  B.
@@
The overall loss is simply an addition of the above two expressions, $\mathcal{L}(\theta)=\mathcal{L}_{CE}(\theta)+\mathcal{L}_{BCE}(\theta)$. This formulation ensures that the model learns to distinguish effectively between real and generated images, as well as classifying them correctly, enhancing its ability to deal with adversarially altered images and to generalize well across different types of image manipulations. More importantly, due to the adversarial training, the learned model has meaningful gradients (i.e., PAG  tsipras2018robustness ,  engstrom2019adversarial ), which are about to play a vital role in our inference process of pushing Consistency generated images to their improved versions. As our training strategy considers both classification and discrimination losses, these gradients carry both improvements geared by class adjustments and a migration from $p_{\theta}$ to $p_{data}$.
In our training process, as described above, we initialize the joint model with a robust RN50 network  madry2018towards , while replacing its fully connected head with a randomly initialized one. As a consequence, we expect this network to inherit it’s original classification robustness, which proves to be valuable for the overall image improvement obtained hereafter. Our training stands on the shoulders of this classifier, as we update the model parameters with the loss described above using only attacked (real and synthesized) images.